what is the hypotenuse of a triangle with the sides of 12 and 12?

Found 2 solutions by Fombitz, Cromlix:
Answer by Fombitz(32388)
Answer by Cromlix(4381)

Hi there,
By Pythagoras Theorem
12^2 + 12^2 = Hypotenuse^2
Hypotenuse = √144 + 144
Hypotenuse = 16.97 (2 decimal places)
Hope this helps :-)